# GPU Cluster Scheduling for Network-Sensitive Deep Learning

### v1

@@ 4.3.1. Preemption priority
As already discussed in Section 3, Tiresias’s 2DAS based priority mechanism is oblivious to network slowdowns. In a large networked cluster, an effective priority mechanism should consider the relative network sensitivity of jobs and the slowdowns they have experienced. Therefore, we propose a network-sensitive priority metric to address this shortcoming. We define the number of iterations completed as $I_{compl}$ and the total expected iterations as $I_{total-expected}$. Our approach necessitates awareness or prediction of the overall expected iteration count. This information can be readily obtained from openly accessible benchmarks like MLperf (mlperf_mlsys20, ). Alternatively, users can estimate the total iteration count needed for convergence through established extrapolation techniques such as (optimus, ). It is important to emphasize that only an approximate value for the total expected iterations suffices.
@@
Next, we define $T_{run}$ as the time spent by the job in the run queue and the total ideal job running time $T_{total-ideal-run}$ as the compute time per iteration of a job multiplied by the total number of expected iterations. The compute time per iteration can be measured by running a single training iteration of the model on a single GPU, where the iteration comprises the forward pass and backward pass (we assume no communication overhead for ideal JCT calculations). We now define work completed $W_{compl}$ and normalized running time $T_{norm}$ as:
@@
4.3.2. Job placement
In a DL cluster, jobs may receive resource offers with varying levels of consolidation, meaning the GPUs offered may be located on a single machine, within a single rack, or spread across the network. Dally employs the well-established delay scheduling strategy to handle these offers. Under this scheme, a job has the option to decline offers that do not meet its consolidation preference as described in Algo. 1. Initially, a job may reject all offers that do not place its GPUs on the same machine until a machine-level delay timer has elapsed while waiting (starving) for resources (line 10 in Algo. 1). Once the machine-level delay timer elapses, the job will commence accepting offers that position all its GPUs within a single rack until the rack-level delay timer has also elapsed (line 18 in Algo. 1). Subsequently, at this point, the job will accept any resource offer without considering consolidation (line 21 in Algo. 1). Note that a job that cannot fit on a single machine will have a machine-level delay timer set to 0, and similarly, a job that cannot fit on a single rack will have both delay timers set to 0.
The underlying principle is to delay the placement of jobs in anticipation of better consolidation options becoming available in the future. This delay-based approach strikes a balance between the high queuing times associated with consolidation and the high communication overhead of non-consolidation. In traditional big data schedulers, these delay timers are typically configured by system administrators and often left at their default values of a few seconds. Based on our empirical observations, we establish a default value of 12 hours for machine-level consolidation and another 12 hours (or 24 hours in total) for rack-level consolidation. Delay timers need to be adjusted based on resource contention in the cluster for optimal delay scheduling. Generally, system administrators adjust timers based on historical contention data of the cluster. But, to eliminate the need for manual timer adjustments, we introduce an ”auto-tuner” that automatically optimizes these delay timers. We elaborate on our auto-tuner in the following section.
4.3.3. Auto-tuner
We introduce an auto-tuner module in Dally aimed at automatically fine-tuning delay timers. This tuner relies on ”moving averages” derived from historical job waiting (starvation) times associated with specific consolidation and GPU demands. The core concept revolves around leveraging historical data to determine an optimal waiting (delay) duration for securing a favorable job consolidation. When a job accepts a resource request, it updates a ”global list” pertaining to the consolidation level (machine or rack) and the job’s GPU demand, with the time it waited for before accepting a resource offer (lines 7 and 15 in Algo. 1). Dally maintains comprehensive lists, tracking job waiting times for each combination of consolidation and GPU demand. In a DL cluster, the GPU demands usually align with multiples of 2, typically ranging from 5 to 10 types. As a result, this approach leads to a manageable 2-20 lists for machine-level and rack-level consolidation paired with GPU demands.
To calculate the moving averages (described in Algo. 2), Dally mandates users to define the size of the wait-time list, denoted as the ”history limit” (sliding window size) for the auto-tuner. The maintenance of the wait-time lists involves discarding values (utilized in average calculation) that exceed the specified history time limit (lines 4 and 10 in Algo. 2). Whenever a job receives a resource offer, the delay timer is calculated by averaging the wait-time list corresponding to the consolidation level (machine or rack) and GPU demand of the job, and then adding two sample standard deviations (similarly estimated using the same wait-time list). This value is then used in determining the optimal duration for a job to wait for a favorable consolidation (lines 13 and 14 in Algo. 2). We present the timeline depicting the adjustment of delay times (rack level) for a typical ¿P95 DDL job in Fig. 5. It is evident from our observations that the auto-tuner modifies delay timers by increasing the value when there is a lower number of racks (and consequently fewer GPUs) in the cluster leading to more resource contention. Over time, the auto-tuner effectively ”learns” the optimal duration to wait as demand delays are updated.
Figure 5. Auto tuning timeline
Note that it is essential to tailor the history limit according to the cluster size. Specifically, in larger clusters, a smaller history time limit is advised. This is simply because more jobs get placed over time in a large cluster compared to a smaller cluster.
@@ 4.3.3. Auto-tuner
4:$T_{starvation}=T_{curr\_time}-T_{last\_assigment}$
5:$g_{d}=\mbox{GPUs demanded by the job}$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
5. EVALUATION
In this section, we assess Dally through extensive simulations driven by large-scale traces, following the previously outlined methodology. It is important to emphasize that our evaluations leverage the most recent generation of network hardware, necessitating the use of simulations as the exclusive option.
5.1. Experimental setup
5.1.1. Workloads
@@
In our experiment, we employ real-world production traces obtained from SenseTime. We randomly select 500 DL/DDL jobs for batch and about 400 DL/DDL jobs for Poisson workloads from the trace, [^3] From the original SenseTime trace, we replace EfficientNet with AlexNet for a high skew workload which utilize the models specified in Table 1. The table encompasses a diverse range of network overheads, providing ample coverage for testing a DDL cluster executing various network-sensitive DNN models. The cluster trace furnishes the compute requirements for individual jobs, including the time per iteration, the necessary number of iterations, and the quantity of GPUs demanded. We assume a data-parallel distributed training paradigm.
5.1.2. Cluster setup
@@
In our experiments, we adopt the configuration of a three-tier network data center. Each machine is equipped with eight GPUs interconnected through NVIDIA NVSwitch. Within each rack, eight machines are connected via the NVIDIA Quantum switch, and the racks across the cluster are interlinked through the NVIDIA Spectrum network switch. To set up our network simulations in ASTRA-sim, we rely on the QoS data advertised by NVIDIA (nvidia-quantum, ; nvidia-spectrum, ). Our experimentation involves manipulating the number of racks, specifically using 2, 4, 8, and 16 racks, to observe the impact of varying compute and network resources.
@@
Figure 6. Batch arrival results: (a) Makespan; (b) Tail queueing delay; (c) Average communication overhead
5.1.3. Baselines
We use the following three baselines to evaluate our scheme:
  1. (1)
@@ 5.1.3. Baselines
  3. (3)
    Dally-manual: This represents a modification of our proposed scheme where delay timers for delay scheduling are manually configured. This aligns with the approach utilized by contemporary state-of-the-art big-data cluster schedulers, such as YARN. Note that the preemption mechanism remains consistent with our proposed scheme Dally with auto-tuning. We use this baseline to demonstrate the effectiveness of our delay timer auto-tuner.
5.1.4. Performance metrics
We define the following metrics to measure performance benefits.
  1. (1)
    Makespan: The time elapsed between arrival of the first job to the completion of the final job in the cluster.
  2. (2)
@@ 5.1.4. Performance metrics
  4. (4)
    Communication overhead: Exposed communication time of a job.
5.2. Results
@@
Figure 7. JCT CDF - Batch arrival
@@
Figure 8. JCT histogram (P95) - Poisson arrival
5.2.1. Makespan
@@
Fig. 6 (a) illustrates the improvement in Makespan. Notably, Dally exhibits an improvement of up to 69% compared to Tiresias and up to 92% compared to Gandiva. This improvement is attributed to the implementation of delay scheduling, effectively relaxing stringent consolidation, thereby decreasing queueing delay and optimizing JCT by judiciously awaiting resource offers with favorable consolidation. A loosening essentially implies a job being positioned in less favorable network tiers, a scenario which may not be feasible in Tiresias. The manual configuration of delay timers in Dally-manual results in sub-optimal waiting, as it cannot dynamically ”tune” its delay timers based on contention. The static nature of the delay time fails to consider GPU demand and resource contention in the cluster, leading to inferior performance compared to Dally, while still outperforming Tiresias and Gandiva due to reduced queuing delays.
5.2.2. Queueing delay
@@
The primary contributor to the significant improvement in Makespan is the reduction in queueing delays experienced within the cluster. Although Dally reports an average queueing delay improvement ranging from 18% to 31%, depending on the cluster size, the cluster’s Makespan is more notably influenced by the ¿P95 and ¿P99 JCTs (i.e. the 95th and 99th complementary percentiles of JCT distribution). Dally demonstrates a remarkable improvement of 58% – 71% in ¿P95 and 67% – 78% in ¿P99 queueing delay compared to Tiresias (Fig. 6 (b)), which results in the substantial improvement in Makespan. The reduction in queueing delay can be attributed to the less stringent consolidation policy of Dally compared to Tiresias, although not as relaxed as Gandiva’s policy. Yet, Dally improves upon Gandiva in tail (¿P95 and ¿P99) queueing delay by as much as 81%. In fact, Gandiva’s network-agnostic scheduling policy leads to very high communication overhead.
5.2.3. Communication overhead

### v2

@@ IV-B1 Preemption Priority
As already discussed in Section III, Tiresias’s 2DAS based priority mechanism is oblivious to network slowdowns. In a large networked cluster, an effective priority mechanism should consider the relative network sensitivity of jobs and the slowdowns they have experienced. Therefore, we propose a network-sensitive priority metric to address this shortcoming. We define the number of iterations completed as $I_{compl}$ and the total expected iterations as $I_{total-expected}$. Our approach necessitates the awareness of the overall expected iteration count which in general is a user input (hyper-parameter). Alternatively, users can estimate the total iteration count needed for convergence through established extrapolation techniques such as [optimus]. It is important to emphasize that only an approximate value for the total expected iterations suffices.
@@
Next, we define $T_{run}$ as the time spent by the job in the run queue and the total ideal job running time $T_{total-ideal-run}$ as the compute time per iteration of a job multiplied by the total number of expected iterations. The compute time per iteration can be measured by running a single training iteration of the model on a single GPU, where the iteration comprises the forward pass and backward pass (we assume no communication latency for ideal JCT calculations). Profilers like DS-analyzer [ds-analyzer] may be used to find the exact compute time of an iteration. We now define work completed $W_{compl}$ and normalized running time $T_{norm}$ as:
@@
IV-B2 Job Placement
@@
[…]
@@
9:  $\mbox{update\_demand\_delay}(machine,t_{starvation},g_{d})$
@@
17:  $\mbox{update\_demand\_delay}(rack,t_{starvation},g_{d})$
@@
In a DL cluster, jobs may receive resource offers with varying levels of consolidation, meaning the GPUs offered may be located on a single machine, within a single rack, or spread across the network. Dally employs the well-established delay scheduling strategy to handle these offers. Under this scheme, a job has the option to decline offers that do not meet its consolidation preference as described in Algo. 1. Initially, a job may reject all offers that do not place its GPUs on the same machine until a machine-level delay timer has elapsed while waiting (starving) for resources (line 10 in Algo. 1). Once the machine-level delay timer elapses, the job will commence accepting offers that position all its GPUs within a single rack until the rack-level delay timer has also elapsed (line 18 in Algo. 1). Subsequently, at this point, the job will accept any resource offer without considering consolidation (line 21 in Algo. 1). Note that a job that cannot fit into a single machine will have a machine-level delay timer set to 0, and similarly, a job that cannot fit into a single rack will have both delay timers set to 0.
The underlying principle is to delay the placement of jobs in anticipation of better consolidation options becoming available in the future. This delay-based approach strikes a balance between the high queuing times associated with consolidation and the high communication latency of non-consolidation. In traditional schedulers, these delay timers are typically configured by system administrators and often left at their default values of a few seconds. Based on our empirical observations, we establish a default value of 12 hours for machine-level consolidation and another 12 hours (or 24 hours in total) for rack-level consolidation. Given the extended duration of DL jobs (days and weeks), it is reasonable to consider delay timers ranging from half a day to a full day, in contrast to the more typical timer values of just a few seconds used in Spark [spark] and YARN [yarn] for shorter (few minutes) jobs. Generally, system administrators adjust timers based on historical resource contention data of the cluster. But, to eliminate the need for manual timer adjustments, we introduce an “auto-tuner” that automatically optimizes these delay timers. Remember that this is an MLaaS provider and will often encounter “new” models.
@@
3:mc_starvation_timesmc\_starvation\_times = $\mbox{get\_delay\_list}(machine,G_{d})$
@@
5:  if $\text{time}>\text{HISTORY\_TIME\_LIMIT}$ then
@@
9:rack_starvation_timesrack\_starvation\_times = $\mbox{get\_delay\_list}(rack,G_{d})$
@@
11:  if $\text{time}>\text{HISTORY\_TIME\_LIMIT}$ then
@@
15:return $\text{average}(\text{mc\_starvation\_times})+2\times\text{standard\_deviation}(\text{mc\_starvation\_times}),$
16:       $\text{average}(\text{rack\_starvation\_times})+2\times\text{standard\_deviation}(\text{rack\_starvation\_times})$
IV-B3 Auto-tuner
We introduce an auto-tuner module in Dally aimed at automatically fine-tuning delay timers. This tuner relies on “moving averages” derived from historical job waiting (starvation) times associated with specific consolidation and GPU demands. The core concept revolves around leveraging historical data to determine an optimal waiting (delay) duration for securing a favorable job consolidation. When a job accepts a resource request, it updates a “global list” pertaining to the consolidation level (machine or rack) and the job’s GPU demand, with the time it waited for before accepting a resource offer (lines 7 and 15 in Algo. 1). Dally maintains comprehensive lists, tracking job waiting times for each combination of consolidation and GPU demand. In a DL cluster, the GPU demands usually align with multiples of 2, typically ranging from 5 to 10 types. As a result, this approach leads to a manageable 2-20 lists for machine-level and rack-level consolidation paired with GPU demands.
To calculate the moving averages (described in Algo. 2), Dally mandates users to define the size of the wait-time list, denoted as the “history limit” (sliding window size) for the auto-tuner. The maintenance of the wait-time lists involves discarding values (utilized in average calculation) that exceed the specified history time limit (lines 4 and 10 in Algo. 2). Whenever a job receives a resource offer, the delay timer is calculated by averaging the wait-time list corresponding to the consolidation level (machine or rack) and GPU demand of the job, and then adding two sample standard deviations [^1] Two standard deviations above mean corresponds to 95% confidence and is commonly made in network system performance evaluation [2-sigma]. (similarly estimated using the same wait-time list). This value is then used in determining the optimal duration for a job to wait for a favorable consolidation (lines 13 and 14 in Algo. 2). We present a timeline depicting the adjustment of delay timers (rack level) for a typical P95 DDL job in Fig. 4. In the figure, the auto-tuner modifies delay timers by increasing or decreasing their values based on resource contention. Clusters with higher resource contention (smaller clusters) have larger delay values. Over time, the auto-tuner effectively “learns” the optimal duration to wait as demand delays are updated.
Figure 4: Auto-tuning timeline for rack-level delay timers.
Note that it is essential to tailor the history limit according to the cluster size. Specifically, in larger clusters, a smaller history time limit is advised. This is simply because more jobs get placed over time in a larger cluster compared to a smaller cluster.
IV-C Simulation Methodology
IV-C1 Simulator design
ArtISt-sim builds upon a prior work, Themis [themis-nsdi], which is a DL cluster scheduler simulator, and ASTRA-sim. The Themis simulator simulates multiple jobs in a data center like a top level cluster scheduler such as Kubernetes [kubernetes]. The choice of Themis is based on its unique capability to simulate DL jobs at an iteration level – a feature not present in other simulators like [lucid-asplos23, chronus-socc21, tiresias]. This granular simulation approach enables us to precisely track “job progression” for preemption and subsequent placement decisions throughout the job’s lifetime. However, as previously discussed, the network slowdown in the original work is encoded statically i.e., hard coded static values for rack and inter-rack placements with no provision to accommodate any other type of inter node network connection. To address this limitation, in ArtISt-sim, we modified Themis to invoke ASTRA-sim for each new job placement which can now be of any given type (e.g., NVSwitch, RoCE, Infiniband, ethernet etc.) as described by the user (refer to Fig. 5).
For executing ASTRA-sim (a single DDL simulator) with specific network placements, ArtISt-sim dynamically generates the necessary input – workload, system, and network topology files – and passes them to ASTRA-sim (downward blue arrow in Fig. 5) to simulate a specific job. These generated files delineate the “portion” of the datacenter network utilized by the worker GPUs. For instance, in Fig. 2, the grey-colored job occupies 8 GPUs across two machines connected via a network link. To simulate ASTRA-sim for such a placement, ArtISt-sim produces the 8-GPU configuration files with a two-tier hierarchy representing the placement. The configuration files specify that the intra-machine GPUs are connected via NVSwitch in a ring topology, and the two machines are linked via a network (Ethernet) switch using a switch topology. Other ASTRA-sim configurations, such as collectives, are also dynamically generated. ASTRA-sim then accurately simulates the specific job for a single DDL iteration based on the placement and relays the communication latency back to ArtISt-sim (upward blue arrow in Fig. 5). This latency is then utilized to simulate the job’s progression (iterations) until it is either interrupted due to preemption or successfully completes its execution.
In Fig 6 we show the timeline of communication latency applied by Themis and ArtISt-sim for 6 jobs. Themis simply applies a fixed latency penalty based on the “network tier” (machine, rack or network) of the jobs placement. However, ArtISt-sim dynamically calculates the latency of each job based on the specific placement at the given time for a high fidelity simulation. In the figure, ArtISt-sim assigns a unique latency for each of the jobs. Note that the compute time per iteration is extracted from the workload trace, and ASTRA-sim serves solely to calculate the network latency for the specific placement.
Figure 5: Simulation design.
Figure 6: Job communication latency timeline.
IV-C2 Simulation calibration and accuracy
To ensure the fidelity of our simulations, we employ a calibration process that incorporates real-world measurements into ArtISt-sim. We execute each model listed in Table I on an 8-GPU machine connected via NVSwitch and calculate the associated network latency. Along with this, we simulate the model with the same network configurations (NVSwitch) in ASTRA-sim. Subsequently, we scale the workload configuration file of ASTRA-sim (for each model) to eliminate the discrepancy between simulated and actual network latency. For each DNN workload, the error in simulation is brought down to be within 1%. These modified workload files are then used for all ASTRA-sim calls. The “calibrated” simulator is now ready to simulate latency arising from rack and network-level placements. Also, for the verification of the various baseline scheduler implementations in ArtISt-sim, we rely on the fact that Themis has already been validated in [themis-nsdi] with real world experiments i.e. each baseline in the simulator was also run in real world setup. Note that only a simulator version of the Tiresias source code is publicly available and no source code of Gandiva is publicly available. For ASTRA-sim, the fidelity has been validated in various prior work such as [rashidi2020astra, won2023astrasim20, rashidi2022themis, Akella22]. Additionally, we faithfully reproduced the design principles of the Tiresias and Gandiva schedulers as outlined in their respective publications.
IV-C3 System Workloads
We employ the cluster job traces made available by SenseTime [sensetime] through the work presented in [lucid-asplos23]. These traces are more recent than those utilized in [philly_trace_analysis] and [weng2022mlaas], and also provide a more realistic representation of a real-world DL cluster with multiple jobs being submitted over an extended period. This trace encompasses individual job arrivals alongside pertinent metadata information.
IV-C4 Implementation
We have implemented ArtISt-sim in about 2000 lines of Java. ArtISt-sim requires the network configuration parameters for intra-machine, intra-rack, and inter-rack interconnects. These parameters encompass the network topology, bandwidth, latency, collective communication algorithm, and other relevant details. Additionally, the user needs to specify the number of machines per rack and the total number of racks in the cluster.
V EXPERIMENTAL METHODOLOGY
This section details the workloads, setup, baselines and metrics used in our evaluation.
V-A Workloads
@@
In our experiment, we employ real-world production traces obtained from SenseTime. We randomly select 500 DL/DDL jobs for batch and about 400 DL/DDL jobs for Poisson workloads from the trace, [^2] From the original SenseTime trace, we replace EfficientNet with AlexNet for a high skew workload which utilize the models specified in Table I. The table encompasses a diverse range of network latency, providing ample coverage for testing a DDL cluster executing various network-sensitive DNN models. The cluster trace furnishes the compute requirements for individual jobs, including the time per iteration, the necessary number of iterations, and the quantity of GPUs demanded. We assume a data-parallel distributed training paradigm. Note that both our batch and Poisson experiments push our system into a congested network regime (a common occurrence in multi-tenant environments [philly_trace_analysis]), which is the scenario of interest.
V-B Cluster Setup
@@
In our experiments, we adopt the configuration of a three-tier network data center. Each machine is equipped with eight GPUs interconnected through NVIDIA NVSwitch. Within each rack, eight machines are connected via the NVIDIA Quantum switch, and the racks across the cluster are interlinked through the NVIDIA Spectrum network switch. To set up our network simulations in ASTRA-sim, we rely on the QoS data advertised by NVIDIA [nvidia-quantum, nvidia-spectrum]. Our experimentation involves manipulating the number of racks, specifically using 2, 4, 8, and 16 racks, to observe the impact of varying compute and network resources.
Figure 7: Batch arrival results – Makespan
@@
Figure 8: Batch arrival results – (a) Tail queueing delay; (b) Average communication latency.
V-C Baselines
In our evaluations, we performed experiments with the following schedulers:
  1. 1.
@@ V-C Baselines
  3. 3.
    Dally-manual: This represents a modification of our proposed scheme where delay timers for delay scheduling are manually configured. This aligns with the approach utilized by contemporary state-of-the-art big-data cluster schedulers, such as YARN.
  4. 4.
    Dally-noWait: Yet another modification of our scheme where the delay timers are set to 0 i.e., the scheduler does not wait for a consolidated placement if none is available.
  5. 5.
    Dally-fullyConsolidated: This final modification of our scheme disregards delay timers and waits as long as needed for the most consolidated placement.
Note that the preemption mechanism remains consistent with our proposed scheme Dally with auto-tuning. We use this baseline to demonstrate the effectiveness of our delay timer auto-tuner.
V-D Performance Metrics
We define the following metrics to measure performance benefits of a job.
  1. 1.
    Makespan: The time elapsed between arrival of the first job to the completion of the final job in the cluster. This is also a proxy for cloud-spend as public cloud bills its tenant based on wall-clock time or makespan occupying its resources. Hence, the monetary expenditure is always a multiple of it.
  2. 2.
@@ V-D Performance Metrics
  4. 4.
    Communication latency: Exposed communication time.
VI EVALUATION
In this section, we assess Dally through extensive simulations driven by large-scale traces, following the outlined design. It is important to emphasize that our evaluations leverage the most recent generation of network hardware, necessitating the use of simulations as the exclusive option.
@@ VI EVALUATION
Figure 9: JCT CDF – Batch arrival.
@@
Figure 10: JCT histogram (P95) – Poisson arrival.
VI-1 Makespan
@@
Fig. 7 illustrates the improvement in makespan. Notably, Dally exhibits an improvement of up to 69% compared to Tiresias and up to 92%, compared to Gandiva. This improvement is attributed to the implementation of delay scheduling, effectively relaxing stringent consolidation, thereby decreasing queueing delay and optimizing JCT by judiciously awaiting resource offers with favorable consolidation. A loosening essentially implies a job being positioned in less favorable network tiers, a scenario which may not be feasible in Tiresias. The manual configuration of delay timers in Dally-manual results in sub-optimal waiting, as it cannot dynamically tune its delay timers based on contention. The static nature of the delay time fails to consider GPU demand and resource contention in the cluster, leading to inferior performance compared to Dally, while still outperforming Tiresias and Gandiva due to reduced queuing delays. Dally-noWait experiences high network latency, leading to poorer performance compared to Tiresias. However, in the case of two racks with limited inter-rack communication, it outperforms Tiresias in terms of makespan, thanks to our novel network-sensitive preemption policy. Dally-fullyConsolidated performs worse than Dally but still surpasses Tiresias, as it does not utilize the delay scheduling policy. These results clearly highlight the effectiveness of both our preemption and delay scheduling strategies.
VI-2 Queuing Delay
@@
The primary contributor to the significant improvement in makespan is the reduction in queueing delays experienced within the cluster. Although Dally reports an average queueing delay improvement ranging from 18% to 31%, depending on the cluster size, the cluster’s makespan is more notably influenced by the  P95 and  P99 JCTs (i.e. the 95th and 99th complementary percentiles of JCT distribution). Dally demonstrates a remarkable improvement of 58% – 71% in  P95 and 67% – 78% in  P99 queuing delay compared to Tiresias (Fig. 8 (a)), which results in the substantial improvement in makespan. The reduction in queuing delay can be attributed to the less stringent consolidation policy of Dally compared to Tiresias, although not as relaxed as Gandiva’s policy. Yet, Dally improves upon Gandiva in tail ( P95 and  P99) queueing delay by as much as 81%. In fact, Gandiva’s network-agnostic scheduling policy leads to very high communication latency.
VI-3 Communication latency